# Hinge electronic structure of strained half-Heuslers

### v1

@@ VI Conclusions
In addition, we have found that uniaxial strain reducing the cubic symmetry does not change the topology of the found hinge modes but can affect their energy. This might provide a controllable platform to experimentally study various problems involving electronic matter in one dimension. One relevant open question is the stability of the hinge modes with respect to possible surface reconstructions, known to be relevant for semiconductors within the same space group [66]. Experimental information of the actual hinges that naturally tend to exist in this broad material class would be of interest. Last, in view of their relatively small bandwith, a second interesting question for future investigation is the fate of the found hinge modes in the presence of Coulomb interactions not considered in our tight-binding Hamiltonian description.
@@ VII Acknowledgements
We thank Ulrike Nitzsche for technical assitance. JIF would like to thank the Alexander von Humboldt Foundation for support. This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-id 390858490)
@@ Appendix A Eight-band tight-binding Hamiltonian
Figure 6: Band structure for LiBiZn. Red, blue and green colors indicate the band stuctures obtained from DFT, full Wannier tight-binding, and truncated Wannier tight-binding Hamiltonian, respectively. Qualitatively similar level of agreements between the DFT and Wannier models is obtained for LiSbZn.
Notice that the calculations in the main text are based on the tight-binding models involving all of the ab initio obtained matrix elements. As explained in the main text, at least six orbitals must be considered to correctly describe the topological properties of the electronic structure of hHs. For simplicity in the DFT downfolding, we keep eight orbitals in the energy windows ($[-6,2]\,$eV). We use the basis $\{|s\uparrow\rangle,|s\downarrow\rangle,|p_{y}\uparrow\rangle,|p_{y}\downarrow\rangle,|p_{z}\uparrow\rangle,|p_{z}\downarrow\rangle,|p_{x}\uparrow\rangle,|p_{x}\downarrow\rangle\}$ where $s$ corresponds to Zn-$4s$ and $p$ to either Bi-$6p$ or Sb-$5p$. As explained in the main text, the lattice Bravais vectors $\vec{a}_{1,2,3}$ read
@@
where $\hat{x},\hat{y},\hat{z}$ are the Cartesian vectors and $a$ is the lattice constant. The reciprocal lattice vectors $\vec{b_{1}},\vec{b_{2}},\vec{b_{3}}$ are expressed as  $\vec{b_{1}}=\frac{2\pi}{a}(\hat{x}-\hat{y}+\hat{z})$,  $\vec{b_{2}}=\frac{2\pi}{a}(\hat{x}+\hat{y}-\hat{z})$,  $\vec{b_{3}}=\frac{2\pi}{a}(-\hat{x}+\hat{y}+\hat{z})$.
@@
The resulting Hamiltonian involves diagonal energies for $s$ ($\varepsilon_{s}$) and $p$ orbitals ($\varepsilon_{p}$), a local spin-orbit coupling $\lambda_{soc}$ in the $p$ shell, hoppings between $s$ orbitals ($t_{s}$), between identical ($t^{(d)}_{p}$) or different ($t^{(od)}_{p}$) $p$ orbitals, and between $s$ and $p$ orbitals ($t_{sp}$) as well as a local hybridization between $s$ and $p$ orbitals ($\nu_{sp}$). The values adopted by these parameters in LiBiZn and LiSbZn are presented in Table 1. The Hamiltonian in momentum space reads
$\begin{split}&\mathcal{H}(\mathbf{k})=\mathcal{H}_{0}+\\ &\big{(}t_{sp}h_{sp,001}+t^{(od)}_{p}h^{(od)}_{p,001}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,001}\big{)}e^{i\frac{(k_{x}+k_{y})}{2}}+\\ &\big{(}t_{sp}h_{sp,010}+t^{(od)}_{p}h_{p,010}^{(od)}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h_{p,010}^{(d)}\big{)}e^{i\frac{(k_{x}+k_{z})}{2}}+\\ &\big{(}t_{sp}h_{sp,100}+t^{(od)}_{p}h^{(od)}_{p,100}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,100}\big{)}e^{i\frac{(k_{y}+k_{z})}{2}}+\\ &\big{(}t^{(od)}_{p}h^{(od)}_{p,1-10}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,001}\big{)}e^{i\frac{(k_{y}-k_{x})}{2}}+\\ &\big{(}t_{p}h^{(od)}_{p,10-1}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h_{p,010}^{(d)}\big{)}e^{i\frac{(k_{z}-k_{x})}{2}}+\\ &\big{(}t^{(od)}_{p}h^{(od)}_{p,01-1}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,100}\big{)}e^{i\frac{(k_{z}-k_{y})}{2}}+\\ &+h.c.\end{split}$ (5)
@@
Appendix B Alternate prism geometry

### v2

@@ VI Conclusions
In addition, we have found that uniaxial strain reducing the cubic symmetry does not change the topology of the found hinge modes but can affect their energy. This might provide a controllable platform to experimentally study various problems involving electronic matter in one dimension. One relevant open question is the stability of the hinge modes with respect to possible surface reconstructions, known to be relevant for semiconductors within the same space group [66]. Experimental information of the actual hinges that naturally tend to exist in this broad material class would be of interest. Last, in view of their relatively small bandwidth, a second interesting question for future investigation is the fate of the found hinge modes in the presence of Coulomb interactions not considered in our tight-binding Hamiltonian description.
@@ VII Acknowledgements
We thank Ulrike Nitzsche for technical assistance. JIF would like to thank the Alexander von Humboldt Foundation for support. This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-id 390858490)
@@ Appendix A Eight-band tight-binding Hamiltonian
Figure 6: Band structure for LiBiZn. Red, blue and green colors indicate the band structures obtained from DFT, full Wannier tight-binding, and truncated Wannier tight-binding Hamiltonian, respectively. Qualitatively similar level of agreements between the DFT and Wannier models is obtained for LiSbZn.
Notice that the calculations in the main text are based on the tight-binding models involving all of the ab initio obtained matrix elements. As explained in the main text, at least six orbitals must be considered to correctly describe the topological properties of the electronic structure of hHs. For simplicity in the DFT downfolding, we keep eight orbitals in the energy windows ([6,2][-6,2]\,eV). We use the basis $\{|s\uparrow\rangle,|s\downarrow\rangle,|p_{y}\uparrow\rangle,|p_{y}\downarrow\rangle,|p_{z}\uparrow\rangle,|p_{z}\downarrow\rangle,|p_{x}\uparrow\rangle,|p_{x}\downarrow\rangle\}$ where $s$ corresponds to Zn-$4s$ and $p$ to either Bi-$6p$ or Sb-$5p$. As explained in the main text, the lattice Bravais vectors $\vec{a}_{1,2,3}$ read
@@
[…]
@@
[…]
$\begin{split}&\mathcal{H}(\mathbf{k})=\mathcal{H}_{0}+\\ &\big(t_{sp}h_{sp,001}+t^{(od)}_{p}h^{(od)}_{p,001}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,001}\big)e^{i\frac{(k_{x}+k_{y})}{2}}+\\ &\big(t_{sp}h_{sp,010}+t^{(od)}_{p}h_{p,010}^{(od)}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h_{p,010}^{(d)}\big)e^{i\frac{(k_{x}+k_{z})}{2}}+\\ &\big(t_{sp}h_{sp,100}+t^{(od)}_{p}h^{(od)}_{p,100}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,100}\big)e^{i\frac{(k_{y}+k_{z})}{2}}+\\ &\big(t^{(od)}_{p}h^{(od)}_{p,1-10}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,001}\big)e^{i\frac{(k_{y}-k_{x})}{2}}+\\ &\big(t^{(od)}_{p}h^{(od)}_{p,10-1}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h_{p,010}^{(d)}\big)e^{i\frac{(k_{z}-k_{x})}{2}}+\\ &\big(t^{(od)}_{p}h^{(od)}_{p,01-1}+t_{s}\mathbf{1}_{s}+t^{(d)}_{p}h^{(d)}_{p,100}\big)e^{i\frac{(k_{z}-k_{y})}{2}}+\\ &+h.c.\end{split}$ (5)
@@
Appendix B Momentum-resolved weight functions in LiBiZn
Figure 7: The LiBiZn weight functions $\overline{w}_{\text{H}}$ and $\overline{w}_{\text{S}}$ are averaged over all states with energies $|\varepsilon|<0.05$ eV, in the case of the prism geometry of Fig. 5. The dark red color denotes $\overline{w}_{\text{H}}$, whereas the light blue color represents $\overline{w}_{\text{S}}$. The vertical jumps are finite-size effects. They correspond to momenta at which the number of states in the energy interval changes.
Here we present additional data on the momentum-resolved, averaged weight functions $\overline{w}_{\text{H}}$ and $\overline{w}_{\text{S}}$, in the case of LiBiZn in the prism geometry of Fig. 5. The results in Fig. 7 show the two weights, averaged over all states in the energy interval [0.05,0.05][-0.05,0.05] eV, as a function of $\vec{k}\cdot\vec{a}_{1}$. While close to the Brillouin zone there is almost no weight on the hinge sites, we observe a peak in $\overline{w}_{\text{H}}$ around $\vec{k}\cdot\vec{a}_{1}=1.25$, corresponding to the light-colored regions close to $\varepsilon=0$ in the bandstructure of Fig. 5(b). At this peak, $\overline{w}_{\text{H}}>0.27$ and $\overline{w}_{\text{S}}>0.4$, indicating that more than two thirds of the probability density is concentrated on the boundary of the prism, with more than one quarter concentrated at the hinge sites.
Appendix C Alternate prism geometry